# Fast Revenue Maximization

### v1

@@ Abstract
We study a data-driven problem pricing problem in which a seller offers a price for a single item based on demand observed at a small finite number of historical prices. Our goal is to derive precise evaluation procedures of the value of the historical information gathered by the seller, along with prescriptions for more efficient price experimentation. Our main methodological result is an exact characterization of the maximin ratio (defined as the worst-case revenue garnered by a seller who only relies on past data divided by the optimal revenue achievable with full knowledge of the distribution of values). This result allows to measure the value of any historical data consisting of prices and corresponding conversion rates. We leverage this central reduction to provide new insights about price experimentation. Motivated by practical constraints that impede the seller from changing prices abruptly, we first illustrate our framework by evaluating the value of local information and show that the mere sign of the gradient may sometimes provide significant information to the seller. We then showcase how our framework can be used to run efficient price experiments. On the one hand, we develop a method to select the next price experiment that the seller should use to maximize the information obtained. On the other hand, we demonstrate that our result allows to considerably reduce the price experimentation needed to reach preset revenue guarantees through dynamic pricing algorithms.
Keywords. pricing, data-driven algorithms, robust pricing, value of data, distributionally robust optimization, conversion rate, limited information.
@@ 1 Introduction and Problem Formulation
In this work, we investigate the monopoly pricing problem in which a decision-maker needs to select a price to sell a single good to a consumer. This operational problem crystallizes the trade-off between volume of sales and profit margin as setting a price too low hurts the margin but increases the demand for the product whereas setting it too high has the reverse effect.
An important practical challenge in pricing is that the distribution of willingness-to-pay of consumers, i.e., the largest price at which a consumer is willing to purchase the good, is unknown to the decision-maker. Therefore, decision-makers need to rely on historical information which commonly includes past prices offered and consumers’ purchase decisions at the offered price. Our goal in this work is to provide a principled framework for two complementary tasks: i.) quantify the value of classical data structures in the form of prices and associated conversion rates, and ii.) perform efficient price experimentation.
@@ 1.1 Problem formulation
We consider the problem of a seller trying to sell one indivisible good to one buyer. We assume that the buyer’s value $v$ is drawn from some unknown distribution $F$ with support included in $[\underline{v},\overline{v}]$, where $0\leq\underline{v}<\overline{v}<\infty$. The seller knows that the distribution of values belongs to some non-parametric class of distributions $\pazocal{C}$. The seller also has information about conversion rates observed at historical prices offered. Formally, let $\mathbf{p}=(p_{i})_{i\in\{0,\ldots N+1\}}$ be a non-decreasing sequence of prices such that $p_{0}=\underline{v}$ and $p_{N+1}=\overline{v}$ and let $\mathbf{q}=(q_{i})_{i\in\{0,\ldots N+1\}}$ be a sequence of historical conversion rates observed such that $q_{0}=1$ and $q_{N+1}=0$. The decision-maker information set is defined as
$\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{i},q_{i});\quad i=0\cdots N+1\}.$
The problem we are interested in is the following: how can the seller leverage the information contained in $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ to maximize her revenue, and how should one guide price experimentation.
Pricing mechanisms and performance. The decision-maker chooses a (potentially) randomized pricing mechanism characterized by the cdf of prices the seller posts. We let P={Ψ in $\pazocal{P}=\{\Psi\mbox{ in }\pazocal{D}_{[\underline{v},\overline{v}]}\}$ be the set of randomized mechanisms, where $\pazocal{D}_{[\underline{v},\overline{v}]}$ is the set of probability measures on $[\underline{v},\overline{v}]$.Then, for every $\Psi\in\pazocal{P}$ and any distribution of consumer value $F$, the expected revenue generated from the randomized mechanism is defined as,
@@
where we let $\mathrm{Rev}\left(p|\overline{F}\right):=p\cdot\overline{F}(p)$. The seller aims to select a mechanism whose revenue is as close as possible to the maximal revenue collected with knowledge of the distribution of values $F$. This revenue is generated by a posted price mechanism (Riley and Zeckhauser, 1983) and we define it as $\mbox{opt}(F):=\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|\overline{F}\right).$ We also define the set of optimal prices as $\mathscr{O}(F)=\arg\max_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|\overline{F}\right)$.
In what follows, we evaluate the performance of a mechanism $\Psi$ against a value distribution $F$ through the ratio which compares the revenue generated by the mechanism and the optimal revenue and is defined as,
@@
Nature’s problem. We evaluate the performance of a given pricing mechanism $\Psi$ by considering its worst-case ratio over a class of admissible value distributions denoted as $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and defined as,
$\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})=\left\{F\in\pazocal{C}\text{ s.t. }\overline{F}(p_{i})=q_{i}\text{ for all }(p_{i},q_{i})\in\mathscr{I}_{\mathbf{p},\mathbf{q}}\right\}.$
Hence, the set $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ contains all value distributions in $\pazocal{C}$, which are consistent with the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ observed by the decision-maker. In turn, we define Nature’s problem as,
@@
Given a pricing mechanism $\Psi$, the value of (1.2) is the worst-case performance of the mechanism over any distribution agreeing with $\mathscr{I}_{\mathbf{p},\mathbf{q}}$. A higher ratio indicates that the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ is highly informative in terms of revenue maximization.
Finally, we note that (1.2) quantifies the performance of a fixed policy. In Section 3, we study the value of certain information sets by computing the best achievable worst-case ratio for any pricing mechanism, which is formally defined as,
@@
In this work, we derive results on the performance of pricing mechanisms for two classes of distributions. The class of General distributions denoted as $\pazocal{G}$ that contains all distributions supported on $[\underline{v},\overline{v}]$. We also consider the subset of $\pazocal{G}$ denoted as $\pazocal{F}$ and which denotes the set of regular distributions, i.e., distributions that admit a density, except potentially at the boundaries of their support and that have a non-decreasing virtual value $v-\overline{F}(v)/f(v)$. We note that the regular class of distributions contains a wide variety of distributions; see, e.g., Ewerhart (2013) for a summary of widely used regular distributions. This class is central to the pricing and mechanism design literature and can be alternatively described as the class of distributions that induce a concave revenue function in the quantity space.
@@ 1.2 Contribution
Performance evaluation for any pricing mechanism. To evaluate the performance of different pricing mechanisms one needs to solve the optimization problem defined in (1.2). Doing so is a priori daunting as (1.2) is an constrained infinite-dimensional optimization problem over the space $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ which is a non-parametric subset of value distributions. The constraints arises in the form of shape constraints (e.g. when $\pazocal{C}=\pazocal{F}$ and distributions have to be regular) or constraints on quantiles of the distribution. Our first main contribution consists in reducing the problem (1.2) to one that is amenable to optimization. Formally, we show the following.
@@
Assume $\pazocal{C}$ is equal to $\pazocal{G}$ or $\pazocal{F}$. Fix $N\geq 1$ and let $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ be an information set including $N$ historical prices and conversions rates. Assume $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty then, for any mechanism $\Psi$ in $\pazocal{P}$,
@@
Theorem 1 reduces the initial infinite dimensional optimization problem in (1.2) to a one-dimensional one for any pricing mechanism. It can be used to evaluate and compare pricing-mechanisms under different assumptions about the information the decision-maker has at hand.
@@
We present the proof of Theorem 1 in Section 2. This result significantly generalizes (Allouah et al., 2023, Theorem 1) across several dimensions. First, our reduction also applies when $\pazocal{C}$ is the set of general distributions and therefore allows to investigate settings where the decision-maker does not impose shape constraints. More importantly, our reduction holds for information sets with arbitrary numbers of historical prices $N$ while the analysis in Allouah et al. (2023) only holds for $N=1$. This considerably extends the possible information sets $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ that we can explore and, as we will see, allows us to derive novel insights on price experimentation. We also note that while this work emphasizes on the ratio objective, our proof readily extends to the absolute regret.
@@
New insights on price experimentation. Theorem 1 provides a principled approach to: $i)$ evaluate historical demand observations, $ii)$ prescribe efficient methods to run price experiments.
@@
First, we illustrate in Section 3.1 how our method can be used to shed light on the value of certain type of price experiments. We consider the practical setting where the seller cannot substantially vary the price offered and therefore observed demand realizations at two neighboring prices. We show that our method provides a rigorous way to assess the value of these local experiments and enable us to precisely quantify the value of the gradient as a function of its magnitude. Quite notably, we show that when the demand is regular, the sign of the gradient may already provide considerable information to the seller.
@@
We then use our framework as a way to develop or enrich algorithms for price experimentation. In Section 3.2, we consider a setting in which the seller observed the demand at a single historical point and then may choose a new price to experiment before committing to a revenue maximizing price. We show that by carefully selecting the second price to experiment with, the seller can significantly increase the worst-case ratio, by $10$ to $20\%$ across all possible demand observed at the first price. In addition, in Section 3.3 we demonstrate that our framework can be used to improve dynamic pricing algorithms. By better quantifying the value of data, one can reduce the extent of price experimentation needed to reach a certain revenue guarantee. We consider the case in which the value distribution is regular (and hence the revenue curve is unimodal), and compare our stopping criterion with the commonly used one when experimenting with the Ternary Search algorithm. Our criterion allows to divide by $2$ to $4$ the number of prices offered to ensure a ratio larger than $99\%$. Hence, our procedure can significantly reduce the cost related to price experimentation.
@@ 1.3 Literature Review
Pricing when facing an unknown value distribution has been first considered in a distributionally robust framework with support information (Bergemann and Schlag, 2008; Eren and Maglaras, 2010) and then analyzed with moment information (Chen et al., 2022). Variants of this problem were analyzed when customer commits to a price scheme on a selling horizon (Caldentey et al., 2016) or when the decision-maker chooses among a broader set of mechanisms (see Carroll (2019); Chen et al. (2023); Wang et al. (2024) and references therein).
@@
The pricing problem has then been studied in the offline data-driven setting where the decision-maker has past data about the value distribution and makes a single pricing decision. In the case where past data takes the form of i.i.d. values sampled from the value distribution, Huang et al. (2015) quantifies the number of samples required to be within a factor $(1-\epsilon)$ from the optimal revenue achievable with full knowledge of the distribution. Fu et al. (2015); Babaioff et al. (2018); Daskalakis and Zampetakis (2020); Allouah et al. (2022) provide performance guarantees when the seller has access to limited number of samples.
@@
The data structure we consider in this work is more closely related to the one studied in the dynamic pricing literature where the decision-maker observes noisy demand observation at an offered price. Kleinberg and Leighton (2003); Besbes and Zeevi (2009); Broder and Rusmevichientong (2012); Keskin and Zeevi (2014); Singla and Wang (2023) study the performance of dynamic pricing policies trading-off between exploration and exploitation in parametric and non-parametric settings. Our work differs from this previous literature along various dimensions: we measure the performance of pricing algorithms with respect to the a single-shot pricing decision (as in Leme et al. (2023b)) as opposed to the commonly studied cumulative regret, hence there is no tradeoff between exploration and exploitation in our work. Furthermore, we derive an exact characterization of the worst-case performance of algorithms that allows to exactly quantify the value of past data. All the previous work focuses on rate-optimal guarantees. Finally, our work assumes that the decision-maker has access to non-noisy demand observation.
Our work is more closely related to Allouah et al. (2023), which provides a tractable procedure to arbitrarily approximate the maximin ratio when $N=1$. Our work extends their results to settings where the decision-maker has access to demand information across several prices. This allows us to derive new insights and prescriptions regarding efficient price experimentation.
@@
Finally, we note that our work broadly relates to Meister and Nietert (2021); Leme et al. (2023a); Okoroafor et al. (2023) which study the sample complexity required to estimate the cumulative distribution functions with binary feedback and to Stoye (2009); Besbes and Mouchtaki (2023); Besbes et al. (2023) which provide tight characterizations of the value of data for other problems.

### v2

@@ Abstract
Problem definition: We study a data-driven pricing problem in which a seller offers a price for a single item based on demand observed at a small number of historical prices. Our goal is to derive precise evaluation procedures of the value of the historical information gathered by the seller, along with prescriptions for more efficient price experimentation. Methodology/results: Our main methodological result is an exact characterization of the maximin ratio (defined as the worst-case revenue garnered by a seller who only relies on past data divided by the optimal revenue achievable with full knowledge of the distribution of values). This result allows to measure the value of any historical data consisting of prices and corresponding conversion rates. We leverage this central reduction to provide new insights about price experimentation. Managerial implications: Motivated by practical constraints that impede the seller from changing prices abruptly, we first illustrate our framework by evaluating the value of local information and show that the mere sign of the gradient of the revenue curve at a single point can provide significant information to the seller. We then showcase how our framework can be used to run efficient price experiments. On the one hand, we develop a method to select the next price experiment that the seller should use to maximize the future robust performance. On the other hand, we demonstrate that our result allows to considerably reduce the number of price experiments needed to reach preset revenue guarantees through dynamic pricing algorithms.
Keywords. pricing, experiment, data-driven algorithms, value of data, distributionally robust optimization, conversion rate, limited information.
@@ 1 Introduction and Problem Formulation
In this work, we investigate the monopoly pricing problem in which a decision-maker needs to select a price to sell a single good to a buyer. This operational problem crystallizes the trade-off between the volume of sales and profit margin as setting a price too low hurts the margin but increases the demand for the product whereas setting it too high has the reverse effect.
An important practical challenge in pricing is that the distribution of value of buyers is unknown to the decision-maker. Therefore, decision-makers need to rely on historical information/transaction data, which commonly includes past prices offered and buyers’ purchase decisions at the offered price. Our goal in the present paper is to provide a principled framework for two complementary tasks: i.) quantify the value of classical data structures in the form of prices and associated conversion rates, and ii.) perform efficient price experimentation.
@@ 1.1 Problem formulation
We consider the problem of a seller trying to sell one indivisible good to one buyer. We assume that the buyer’s value $v$ is drawn from some unknown distribution $F$ represented by its cumulative distribution function (cdf) and with support included in $[\underline{v},\overline{v}]$, where $0\leq\underline{v}<\overline{v}<\infty$. The seller knows that the distribution of values belongs to some non-parametric class of distributions $\pazocal{C}$. The seller also has information about conversion rates observed at historical prices offered. Formally, let $\mathbf{p}=(p_{i})_{i\in\{0,\ldots N+1\}}$ be a non-decreasing sequence of prices such that $p_{0}=\underline{v}$ and $p_{N+1}=\overline{v}$ and let $\mathbf{q}=(q_{i})_{i\in\{0,\ldots N+1\}}$ be a sequence of historical conversion rates observed such that $q_{0}=1$ and $q_{N+1}=0$. The decision-maker information set is defined as
$\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{i},q_{i});\;i\in\{0,\ldots,N+1\}\}.$
The problem we are interested in is the following: how can the seller leverage the information contained in $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ to “robustly” maximize her revenue, and how should one guide price experimentation, given that such information has already been collected.
Pricing mechanisms and performance. In this problem, the decision-maker chooses a (potentially) randomized pricing mechanism characterized by the cdf of prices that the seller posts. Formally, we let $\pazocal{P}=\{\Psi\in\pazocal{D}_{[\underline{v},\overline{v}]}\}$ be the set of randomized mechanisms, where $\pazocal{D}_{[\underline{v},\overline{v}]}$ is the set of probability measures on $[\underline{v},\overline{v}]$. Then, for every $\Psi\in\pazocal{P}$ and any distribution of consumer value $F$, the expected revenue generated from the randomized mechanism is defined as,
@@
where we let $\mathrm{Rev}\left(p|F\right)=p\cdot\mathbb{P}_{v\sim F}\left(v\geq p\right).$ To ease notation, we define the complementary cumulative distribution (ccdf) $\overline{F}(p)=1-F(p)$ for every $p\in\mathbb{R}$ and let $\mathbb{P}_{v\sim F}\left(v\geq p\right)=\overline{F}(p-)$. This notation is consistent with the fact that both $\mathbb{P}_{v\sim F}\left(v\geq p\right)$ and $\overline{F}(p-)$ corresponds to the right limit of $\overline{F}$ at $p$.
In an ideal scenario where the seller has full knowledge of the distribution of values $F$, the optimal revenue is obtained by a posted price mechanism (Riley and Zeckhauser, 1983) and we denote it by $\mbox{opt}(F):=\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|F\right).$ We also define the set of optimal prices as $\mathscr{O}(F)=\arg\max_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|F\right)$. In this works, we aim at understanding the revenue gap incurred by the seller due to not knowing the actual distribution. Given a mechanism $\Psi\in\pazocal{P}$ and a value distribution $F$, we define this gap through the following ratio,
@@
Performance for a given information set. Our goal is to understand the spectrum of achievable performances for various information sets $\mathscr{I}_{\mathbf{p},\mathbf{q}}$. In particular, we assume that the seller is given the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$. We then define the performance of a pricing mechanism $\Psi\in\pazocal{P}$ as the worst-case ratio, where the worst-case is taken over a class of admissible value distributions denoted as $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and defined as,
$\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})=\left\{F\in\pazocal{C}\text{ s.t. }\overline{F}(p_{i}-)=q_{i}\text{ for all }(p_{i},q_{i})\in\mathscr{I}_{\mathbf{p},\mathbf{q}}\right\}.$
We note that the set $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ contains all value distributions in $\pazocal{C}$ which are consistent with the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ observed by the decision-maker.
Consequently, the worst-case performance of any pricing mechanism $\Psi\in\pazocal{P}$ given $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ is the value of the following optimization problem.
@@
In what follows, we will refer to (1.2) as Nature’s problem. We note that a higher value of this worst-case ratio indicates that the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ is highly informative in terms of revenue maximization and that the policy $\Psi\in\pazocal{P}$ used by the decision-maker effectively takes advantage of the information in $\mathscr{I}_{\mathbf{p},\mathbf{q}}$.
While (1.2) quantifies the performance of a fixed policy, we study in Section 4 the intrinsic value of certain information sets by computing the best achievable worst-case ratio across all pricing mechanisms. This quantity is formally defined as,
@@
Distribution classes. In this work, we derive results on the performance of pricing mechanisms for two classes of distributions. The class of general distributions denoted as $\pazocal{G}$ that contains all distributions supported on $[\underline{v},\overline{v}]$. We also consider the subset of $\pazocal{G}$ denoted as $\pazocal{F}$ and which denotes the set of regular distributions, i.e., distributions that admit a density, except potentially at the upper boundary of their support and that have a non-decreasing virtual value $v-\frac{1-F(v)}{f(v)}$. We note that the class of regular distributions contains a wide variety of distributions; see, e.g., Ewerhart (2013). This class is central to the pricing and mechanism design literature and can be alternatively described as the class of distributions that induce a concave revenue function in the quantity space.
@@ 1.2 Contribution
Performance evaluation for any pricing mechanism. To evaluate the performance of a given pricing mechanism, one needs to solve the optimization problem defined in (1.2). Doing so is a priori daunting as (1.2) is a constrained infinite-dimensional optimization problem over the space $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ which is a non-parametric subset of value distributions. The constraints arise in the form of shape constraints (e.g., when $\pazocal{C}=\pazocal{F}$, distributions have to be regular) or constraints on quantiles of the distribution. Our first main contribution consists in reducing Problem (1.2) to one that is amenable to optimization. Formally, we show the following.
@@
Assume $\pazocal{C}$ is equal to $\pazocal{G}$ or $\pazocal{F}$. Fix $N\geq 1$ and let $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ be an information set including $N$ historical prices and conversion rates. Assume $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty then, for any mechanism $\Psi$ in $\pazocal{P}$,
@@
For any pricing mechanism, Theorem 1 reduces the initial infinite-dimensional optimization problem in (1.2) whose value corresponds to the robust performance of the pricing mechanism, to a one-dimensional problem. This is a critical reduction that can then be used to evaluate and compare pricing-mechanisms under different assumptions about the information the decision-maker has at hand. It also enables to approximate Problem (Maximin) efficiently and derive near-optimal pricing mechanisms for general historical data structures.
@@
We present the proof of Theorem 1 in Section 2. This result significantly generalizes (Allouah et al., 2023, Theorem 1) across several dimensions. First, our reduction also applies when $\pazocal{C}$ is the set of general distributions and therefore allows to investigate settings where the decision-maker does not impose shape constraints. More importantly, our reduction holds for information sets with an arbitrary number of historical prices $N$, while the analysis in Allouah et al. (2023) only holds for $N=1$. This considerably extends the possible information sets $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ that we can investigate and, as we will see, opens up the possibility to explore prescriptions for price experimentation. In turn, it will also allow us to derive novel insights on price experimentation. We also note that while this work emphasizes the ratio objective, our proof is general and readily applies to the absolute regret objective.
@@
New insights on price experimentation. Theorem 1 provides a principled approach to quantify the value of historical demand observations.
@@
We illustrate in Section 4.1 how our method can be used to shed light on the value of certain types of price experiments. We consider the practical setting in which the seller cannot substantially vary the price offered and observe demand realizations at two neighboring prices. We show that our method provides a rigorous way to assess the value of these local experiments and enables to precisely quantify the value of knowing the gradient of the revenue curve at a single price as a function of its magnitude. Quite notably, we show that when the demand is regular, the mere sign of the gradient may already provide considerable information to the seller. More importantly, our procedure allows to perform a rigorous cost-benefit analysis by identifying scenarios in which the sign of the gradient (which can be estimated with cheaper price experiments) already provides sufficient information for better performances, and scenarios for which the numerical value of the gradient (which requires longer and more expensive experiments to be estimated) is important to pin down.
@@
We then use our framework as a way to develop or enrich algorithms for price experimentation. In Section 4.2, we consider a setting in which the seller observed the demand at a single historical point and then may select a new price to experiment with before committing to a future price. We show that by judiciously selecting the second price to experiment with, the seller can significantly increase the worst-case ratio, by $10$ to $20\%$ across all possible demand observed at the first price.
In Section 4.3 we demonstrate that our framework can be used to improve dynamic pricing algorithms. We illustrate that by better quantifying the value of data, one can reduce the extent of price experimentation needed to reach a certain revenue guarantee. We consider the case in which the value distribution is regular (and hence the revenue curve is unimodal), and compare our stopping criterion with the commonly used one when experimenting with the Ternary search algorithm. For example, our refined stopping criterion allows to divide by $2$ to $5$ the number of prices offered to ensure a ratio larger than $99\%$.
@@ 1.3 Literature Review
Pricing when facing an unknown value distribution has been considered in a distributionally robust framework with support information (Bergemann and Schlag, 2008; Eren and Maglaras, 2010), with moment information (Azar and Micali, 2013; Carrasco et al., 2018; Chen et al., 2022; Chen, Hu and Wang, 2023; Wang et al., 2024) and with knowledge of quantiles (Azar et al., 2013; Allouah et al., 2023). Variants of this problem were also analyzed in the presence of strategic customers (Caldentey et al., 2016).
@@
The pricing problem has also been studied in the offline data-driven setting where the decision-maker has past data about the value distribution and makes a single pricing decision. In the case where past data takes the form of i.i.d. values sampled from the value distribution, Huang et al. (2015) quantifies the number of samples required to be within a factor $(1-\epsilon)$ from the optimal revenue achievable with full knowledge of the distribution. Fu et al. (2015); Babaioff et al. (2018); Daskalakis and Zampetakis (2020); Allouah et al. (2022) provide performance guarantees when the seller has access to a small number of samples. Chen, Cire, Hu and Lagzi (2023) investigate the performance of data-driven pricing algorithms with transactional data and Biggs (2022) proposes pricing policies based on surrogate loss functions to incorporate contextual information.
@@
In the online setting, Aviv and Pazgal (2002); Araman and Caldentey (2009); Farias and Van Roy (2010) examine dynamic pricing within a Bayesian framework, in which the decision-maker possesses prior knowledge about the parameters of the value distribution. The data structure we consider in this work is related to the one studied in the dynamic pricing and learning literature where the decision-maker observes noisy demand observation at an offered price. Kleinberg and Leighton (2003); Besbes and Zeevi (2009); Broder and Rusmevichientong (2012); Keskin and Zeevi (2014); Singla and Wang (2023) study the performance of dynamic pricing policies trading-off between exploration and exploitation in parametric and non-parametric settings. (Broder, 2011, Chapter 3);Cheung et al. (2017); Chen et al. (2020); Perakis and Singhvi (2023) develop and analyze data-driven pricing policies with limited price changes. Bu et al. (2022) investigates a dynamic pricing problem with linear demand, in which the decision-maker uses both offline and online data. They analyze the value of historical (offline) data by quantifying the achievable regret rate as a function of three statistics of the offline transactional data (size, location, and dispersion). Our work differs from this previous literature along various dimensions: we measure the performance of pricing algorithms with respect to a single-shot pricing decision as opposed to the commonly studied cumulative regret, hence there is no tradeoff between exploration and exploitation in our work. Furthermore, we focus on the setting where the decision-maker observes the exact conversion rate at the offered prices and we derive an exact characterization of the worst-case performance of algorithms that allows to exactly quantify the value of past data. In contrast to Bu et al. (2022), our quantification of the value of data is a function of the whole dataset and therefore enables to provide data-dependent insights. We believe that our results and insights may motivate new algorithmic principles for dynamic pricing policies.
Our work relates to Leme et al. (2023b), which compares the informational value of transactional data and actual observations of each buyer’s value. They show that, asymptotically, observing buyer’s value provides significantly more information for revenue maximization. Our work differs by assuming that the decision-maker has a perfect estimate of the conversion rate at a few number of historical prices and characterizes the exact value of such historical data. It is more closely related to Allouah et al. (2023), which provides a tractable procedure to arbitrarily approximate the maximin ratio when the number of conversion rates $N=1$. Our work extends their results to settings in which the decision-maker has access to demand information across several prices. This allows us to derive new insights and prescriptions regarding efficient price experimentation. In parallel developments, that appeared after we posted an initial draft of this paper, Daei Naby et al. (2024) solve the problem when $N=2$ and when the distribution is assumed to be regular. They provide a closed-form characterization of the optimal deterministic price experiment in the same setting as the one we numerically investigate in Section 4.2. Their focus on deterministic prices, closed forms, and associated insights, complements the understanding of this fundamental problem. (Eren and Maglaras, 2010, Section 4) also analyzes the pricing problem with limited experimentation and characterizes the worst-case ratio for general distributions with information about $N$ conversion rates under the assumption that the value distribution is discrete. Our work generalizes this analysis to arbitrary distributions (both continuous and discrete) and provides a different and unified proof which applies not only to general but also to regular distributions.
@@
Finally, we note that our work broadly relates Chehrazi and Weber (2010) which considers a general robust decision problem with limited measurements and applies their methods in the context of optimal debt-settlement, and to Meister and Nietert (2021); Leme et al. (2023a); Okoroafor et al. (2023) which study the sample complexity required to estimate the cumulative distribution functions with binary feedback. We also contribute to the broad literature which aims at deriving an exact characterization of the value of data in the small to moderate data regime (Schlag, 2006; Stoye, 2009; Besbes and Mouchtaki, 2023; Besbes et al., 2023).